OpenStudy (anonymous):

Which system has (-2, -2) as the solution? A) x - y = -2 3x - y = 0 Which system has (1, 3) as the solution? B) x - 4y = -8 2x - 3y = -16 Which system has infinite solutions? C) x + y = -4 -3x + 2y = 2 Which system has (-8, 0) as the solution? D) 4x + 6y = -12 2x + 3y = 9 Which system has no solutions? E) x + y = -1 4x + 4y = -4

OpenStudy (ranga):

Just plug in x and y values and see if the equation is still true or not.

OpenStudy (anonymous):

oh ok thank you.

OpenStudy (ranga):

you are welcome. For A, B and D you can plug in and know which one is correct. I will do just A) x = -2, y = -2. Put it in x - y = -2 -2 - (-2) = -2 + 2 = 0 Left side is 0 but the right side is 2. So this is not the answer. The next one is 3x - y = 0 x = -2, y = -2 3(-2) - (-2) = -6 + 2 = -4. But right side is 0. SO it is not a solution for this one either.

OpenStudy (anonymous):

thank you I got them all right :)

OpenStudy (ranga):

cool.